OpenStudy (anonymous):

HELP ME PLEASE GOLD MEDLE AND FANN I BEGG YOU PLEASE PLEASE I WILL DO ANYTHING FOR U PLEASEEEEEEEEEEEEEEEEE

OpenStudy (anonymous):

\[\frac{ 1 }{ 4 }+\frac{ 2 }{ 7 }\]

OpenStudy (anonymous):

okay first find a common denominator

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

Okay so what is the common denominator

OpenStudy (anonymous):

s1/4 + 2/7 = ((1 × 7) + (2 × 4)) / (4 × 7) = (7 + 8) / 28 = 15/28 = 15/28

OpenStudy (anonymous):

what with the Question marK?

OpenStudy (anonymous):

i could not tell you

OpenStudy (anonymous):

they happen some times

OpenStudy (anonymous):

this site helps me a lot http://www.calculatorsoup.com/calculators/math/mixednumbers.php

OpenStudy (anonymous):

What was the common denominator you got @girlandboysyell!!